

********************************************************************
UARS WEEKLY STATUS REPORT - 10 JULY 1998
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, collecting solar data on all
available orbital tracking periods and stellar calibration observations,
whenever possible.
***********
SUSIM
***********
SUSIM was turned on Friday, 3 July 1998, and, since then, has successfully
gathered its normal solar, offset, and occultation scan data without apparent
problems.
***********
PEM
***********
PEM was not active.
GSFC has proposed an OBC operation plan for the activation of PEM boom
instruments during the daylight portions of every orbit. This OBC operation
plan is accepted by PEM. PEM has proposed a series of test procedures for
activation and is waiting response from GSFC. PEM is also waiting for GSFC
scheduling and implementation of the OBC plan and its operation test procedures.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
MLS has been operated in a very low power short duration diagnostic mode to
obtain engineering data on the scan mechanism. Additional tests are being
conducted to determine the optimal operating mode for the upcoming south
viewing observation period.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no other problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
31 days HAL (4.80) PROD (L1) (1870R-570C-8F=1292L)
55 days HAL (4.80) PROD (L3) (1870R-361C=1509L)
21 days SOL (4.61) PROD (L2-L3) (2349R-708C=1641L)
2 days SUS (7.0) Eng (5R-2C=3L)
The CDHF successfully completed database disk maintenace through the July 3
holiday weekend. The CDHF wished to decrease the impact to CDHF users during
prime time .
CDHF pending software upgrades: IDL 5.1, OpenIngres, Fortran V6.5.
An installation and test area for the OpenIngres 1.1 Upgrade has been set
up.
***********
FOT
***********
REPORTING PERIOD
This report covers 27 June 1998 (Orbit 37113, GMT Day 178) through 03
July 1998 (Orbit 37216, GMT Day 184).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing)
and is performing nominally under two-battery operations. The instrument
operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
ACRIM 178/2235 37126 Off
SUSIM 178/2235 37126 Off
PEM 178/2235 37126 AXIS off
WINDII 178/2235 37126 Off
MLS 179/0129 37128 On
MLS 180/1700 37152 Off
MLS 183/2124 37200 On/Off test
ACRIM 184/0008 37202 On
HRDI 184/0152 37203 On
SUSIM 184/1056 37209 On
The beta angle reached a maximum angle of 33.8 degrees on 30 June 1998
(DOY 181). The beta angle is now decreasing toward a minimum angle of
0.0 degrees on 15 July 1998 (DOY 196).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery
1 remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 24.8 V end-of-night (EON) load bus voltage.
Battery temperatures are stable with a temperature delta between Battery
2 and Battery 3 of 1.0 degrees C. Additional battery performance data for
each battery on 27 June 1998 and 03 July 1998 are provided in Appendix D.
The clock error ranged from -11.6 to 11.1 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and Out-
Of-Limits occurrences are listed in Appendix B.
% ALERTS = none
% OUT-OF-LIMITS = 10
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
% ACRIM 02
% FOT 05
% HRDI 01
% MLS 03
% PEM 01
% SOLSTICE 01
% SUSIM 02
% WINDII 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
% FOT 06
One (1) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 98/146 - 2287.496748
Transponder B: 98/147 - 2287.499000
GROUND SYSTEM OPERATION
Three (3) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 98/183 show the observatory data loss to be
20 hours, 44 minutes, 52.500 seconds (an increase of 14.112 seconds since
last report period). This is a 0.0349 percent data loss which equals a
99.9651 percent data capture for the mission.
The increase of 14.112 seconds occurred over two days:
7 12.096 seconds on Day 180 due to insufficient overlap, and
7 2.016 seconds on Day 182 due to insufficient overlap.
Testing the command side of the I/P net has been scheduled for 7-9 July
1998.
Installed and tested FOT level-0 processing software (UGPL, etc.) on the
UMPGAM computer system. This system can now be used in lieu of or in
parallel with the CDHF and/or UMPGAL systems for processing of level-0
telemetry data.
New AP ground system software testing continued. Several development disk
packs were loaded, but none successfully completed testing. Installation
of the new ground system software is a prerequisite to the uplink of the
new OBC power monitor (PMON) flight software, for which the FOT has
successfully completed all development and testing.
OTHER SIGNIFICANT EVENTS
Planning continues for the UARS orbit adjust, scheduled for 12 August 1998,
and the engineering thruster tests scheduled for 28 July 1998.
FOT continues to improve a power profile for UARS based on the current
status of the Modular Power System.
FOT continues to support the new UARS ground system re-engineering effort.
MEETINGS
A meeting was held on 29 June 1998 to discuss B/D attitude thruster
hydrazine line refill testing. A technique was proposed (using OBC
predefined command blocks) to allow latch valve command spacing down to the
1 msec level for testing purposes. This testing is planned to be performed
in conjunction with the UARS Test and Training Simulator (UTTS).
No UARS Battery telecon was held this week. The next telecon has not been
scheduled.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver is currently scheduled for 15 July 1998.
The UARS orbit altitude is dropping slowly and a drag makeup orbit adjust
is targeted for 12 August 1998.
PERSONNEL CHANGES
None
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 03 July 1998
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
37124 178/1914 98-219 OBC Clk Adj fr 0.00 to 20.6 msec/day
37126 178/2235 98-220 ACRIM Power OFF
37126 178/2235 98-221 PEM Power OFF AXIS
37126 178/2235 98-222 SUSIM Power OFF
37126 178/2235 98-223 WINDII Power OFF
37128 179/0129 98-224 MLS Power ON
37149 180/1158 98-225 OBC Clk Adj fr 20.6 to 0.00 msec/day
37151 180/1407 98-226 OBC Loaded RTS 12-SUSIM Heater Cycle
37152 180/1700 98-227 MLS Power OFF
37186 182/2252 98-228 OBC Clk Adj fr 0.00 to 20.6 msec/day
37197 183/2124 98-229 MLS Test
37202 184/0008 98-230 ACRIM Power ON
37203 184/0152 98-231 HRDI Power ON
37205 184/0603 98-232 SOLSTICE Reboot
37209 184/1056 98-233 SUSIM Power ON
37214 184/1941 98-234 OBC Clk Adj fr 20.6 to 0.00 msec/day
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 03 July 1998
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
QTSISAIF Y-HI SSPP Baseplate by SISA I/F 37129-141,203,204
QTSSPPMX Y-HI SSPP -X Thermostat 37130-171,187-200
HRPSEQBIMON Y-HI HRDI PSE QB Current Mon. 37130-132,145,152
SORLBXTMPR Y-LO Logic_Box Temperature 37132-141,156-171,
37187-37200
QTSISAIF Y-LO SSPP Baseplate by SISA I/F 37135-143,172-186
QTSISAIF R-LO SSPP Baseplate by SISA I/F 37144-156
QTACRIMIF Y-LO SSPP Temp at ACRIM I/F 37156-171
QTSISAIF R-HI SSPP Baseplate by SISA I/F 37156-171,187-200
SORLBXTMPR R-LO Logic_Box Temperature 37156
HRPTSPBIMON Y-HI PTS PB Current Monitor 37204
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
None this report period
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 03 July 1998
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
37124 178/19:20 (27 Jun 98) from 00.0 to 20.6 msec/day
37149 180/12:04 (29 Jun 98) from 20.6 to 00.0 msec/day
37186 182/23:01 (01 Jul 98) from 00.0 to 20.6 msec/day
37214 184/19:57 (03 Jul 98) from 20.6 to 00.0 msec/day
_______________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 03 July 1998
27 Jun 1998 - GMT Day 178 Beta = 31.4 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -2.42/-3.68 +0.0/ +0.0 18.2 0.000 0.0
2 +22.4/ -22.4 +7.09/+5.47 +30.4/-16.0 24.5 1.022 23.1
3 +39.2/ -11.2 +5.79/+4.51 +29.6/-16.0 24.5 1.024 22.5
03 Jul 1998 - GMT Day 184 Beta = 32.5 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -2.42/-3.68 +0.0/ +0.0 18.2 0.000 0.0
2 +11.2/ -50.4 +6.44/+4.83 +29.6/-15.2 24.8 1.021 22.5
3 +39.2/ -11.2 +5.47/+3.87 +28.8/-16.0 24.8 1.024 22.0
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes
APPENDIX E
UARS ANOMALY REPORTS
Week ending 03 July 1998
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-039 FOT 37124 178/19:14 TAC 8 Crash
98-028 FOT 37139 179/20:17 14170 Generic Late Acquisition
98-040 FOT 37142 180/00:18 TAC 8 Crash
98-041 FOT 37175 182/05:58 TAC 8 Crash
98-042 FOT 37177 182/09:08 20031 Late Acquisition
98-043 FOT 37189 183/03:29 TAC 2 Crash
98-044 FOT 37196 183/15:38 14170 Generic Late Acquisition
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-028 FOT See list June 1998 14170 Generic Late Acquisition
98-039 FOT 37124 178/19:14 TAC 8 Crash
98-040 FOT 37142 180/00:18 TAC 8 Crash
98-041 FOT 37175 182/05:58 TAC 8 Crash
98-042 FOT 37177 182/09:08 20031 Late Acquisition
98-043 FOT 37189 183/03:29 TAC 2 Crash
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-044 FOT See list July 1998 14170 Generic Late Acquisition